OpenStudy (anonymous):

The perimeter of a rectangle is 72 centimeters. The base is 2 times the height. What are the dimensions of the rectangle?

OpenStudy (imstuck):

the perimeter of a rectangle is P=2L + 2W, right? Or in our case, P=2B + 2H. We have a perimeter length of 72, so 72= 2B + 2H. We also know that the base is twice the height so height is H and base is 2H. Now we have an equation in terms of H only.

OpenStudy (imstuck):

72=2(H)+2(2H) 72=2H + 4H 72=6H 12=H